OpenStudy (mendicant_bias):

I'm reviewing solving linear, first order ODE's using an integrating factor at the moment and am running into trouble doing the problem in full; posted below shortly.

OpenStudy (mendicant_bias):

\[xy'-y=x^3+1\]

OpenStudy (mendicant_bias):

The trouble that I'm having is that I either have forgotten how to use an integrating factor properly (entirely possible) or that I'm making an error in computing it. Workings so far posted below; I understand that these steps can be skipped, but I'd rather not just use the formula and work through it when practicing.

OpenStudy (mendicant_bias):

First, getting the ODE in the appropriate form to use an integrating factor: \[\frac{dy}{dx}-\frac{y}{x}=\frac{x^3+1}{x}\]

OpenStudy (mendicant_bias):

From here, multiplying both sides by our mystery integrating factor, mu: \[\mu \frac{dy}{dx}-\mu \frac{y}{x}=\mu \frac{x^3+1}{x}\]

OpenStudy (mendicant_bias):

\[\mu ' (x)=\mu (x) P(x)\] by definition for this to be true; solving for mu in the typical way, where it yields: \[\mu = e^{\int\limits_{}^{}P(x)dx}\]

OpenStudy (mendicant_bias):

\[\mu = e^{\int\limits_{}^{}-\frac{1}{x}dx}=e^{-\ln(x)}=e^{\ln(x^{-1})}=\frac{1}{x}\]

OpenStudy (mendicant_bias):

Here's the problem, though, when I take this value and plug it in and differentiate mu*y, I don't get my original problem.

OpenStudy (bohotness):

oh okay

OpenStudy (bohotness):

so that means you need to plug in something else

OpenStudy (mendicant_bias):

Because, for the problem to work, \[\frac{d}{dx}(\mu y)=\mu \frac{dy}{dx}+\mu' y\] Well...like what? Where did I make a mistake, and how?

OpenStudy (bohotness):

okay ley me check it okay

OpenStudy (mendicant_bias):

\[\frac{d}{dx}\bigg(\frac{y}{x}\bigg)=\frac{dy}{dx}\frac{1}{x}-\frac{1}{x^2}y\] I'm getting the original thing in form, divided by a single x.

OpenStudy (bohotness):

okay

OpenStudy (bohotness):

how did you get you answer so i may known where you went wrong

OpenStudy (mendicant_bias):

Well, I got mu doing the typical integrating factor process, outlined above (does that make sense?) and then I differentiated mu times y, which should yield the original equation in that form where y' has a coefficient of mu. The original equation, when put in form, should just be the result of the product rule for differentiation, between mu and y.

OpenStudy (bohotness):

hmmm

OpenStudy (bohotness):

do you have answer chioceS?

OpenStudy (displayerror):

Shouldn't it look like this after multiplying by your integrating factor? \[\frac{d}{dx}(\frac{1}{x}y)= \frac{1}{x} \frac{x^3 +1}{x}\]

OpenStudy (mendicant_bias):

It's not multiple choice, I'm given the answer but regardless, this is prior to the final answer; I need to clear up this conceptual misunderstanding prior to anything else, because either my integrating factor is wrong or the process itself is wrong; Yup, that's what the RHS should look like, I'm confused about the LHS

OpenStudy (mendicant_bias):

I feel like this is something really simple that I'm missing.

OpenStudy (bohotness):

what do you feel like thats missing

OpenStudy (mendicant_bias):

Lol, I don't know. @Kainui , could you help me out on this?

OpenStudy (bohotness):

displayerror you right:3

OpenStudy (mendicant_bias):

Nevermind, I'm an idiot, I think I got it.

OpenStudy (bohotness):

:D

OpenStudy (bohotness):

your not a idiot

OpenStudy (mendicant_bias):

Just forgot how to properly use an integrating factor for a moment, either way, still need to finish the problem in its entirety.

OpenStudy (bohotness):

your smart

OpenStudy (bohotness):

:D yes

OpenStudy (mendicant_bias):

\[\frac{d}{dx}\bigg(\mu (x)y(x)\bigg)=\mu(x)f(x); \ \ \ y(x)=\frac{\int\limits_{}^{}\mu(x)f(x)}{\mu(x)}\]

OpenStudy (bohotness):

:D

OpenStudy (anonymous):

\[\frac{ dy }{ dx }+Py=Q,\] where P and Q are functions of x has an integrating factor \[I.F.=e ^{\int\limits P dx}\] and complete sol. is \[y*I.F=\int\limits Q*I.F. dx+c\]

OpenStudy (mendicant_bias):

\[y=\frac{\int\limits_{}^{}\frac{x^3+1}{x^2}}{\frac{1}{x}}\]

OpenStudy (bohotness):